# BRIGHT: A globally distributed multimodal building damage assessment dataset with very-high-resolution for all-weather disaster response

### v1

@@ Abstract
Disaster events occur around the world and cause significant damage to human life and property. Earth observation (EO) data enables rapid and comprehensive building damage assessment, an essential capability in the aftermath of a disaster to reduce human casualties and to inform disaster relief efforts. Recent research focuses on the development of artificial intelligence (AI) models to achieve accurate mapping of unseen disaster events, mostly using optical EO data. However, solutions based on optical data are limited to clear skies and daylight hours, preventing a prompt response to disasters. Integrating multimodal EO data, particularly the combination of optical and synthetic aperture radar (SAR) imagery, makes it possible to provide all-weather, day-and-night disaster responses. Despite this potential, the development of robust multimodal AI models has been constrained by the lack of suitable benchmark datasets. In this paper, we present a Building damage assessment dataset using veRy-hIGH-resoluTion optical and SAR imagery (BRIGHT) to support AI-based all-weather disaster response. To the best of our knowledge, BRIGHT is the first open-access, globally distributed, event-diverse multimodal dataset specifically curated to support AI-based disaster response. It covers five types of natural disasters and two types of man-made disasters across 12 regions worldwide, with a particular focus on developing countries where external assistance is most needed. The optical and SAR imagery in BRIGHT, with a spatial resolution between 0.3 and 1 meters, provides detailed representations of individual buildings, making it ideal for precise damage assessment. In our experiments, we have tested seven advanced AI models trained with our BRIGHT to validate the transferability and robustness. The corresponding experimental results can also serve as a baseline for future research, providing perspectives and inspiration for the design of custom models. The dataset, along with code and trained models, is available at https://github.com/ChenHongruixuan/BRIGHT, and will be updated as new disaster data become available. BRIGHT also serves as the official dataset for the 2025 IEEE GRSS Data Fusion Contest Track II. We hope that the new dataset will further the use of AI-driven EO methods in support of people in disaster-affected areas.
@@
A disaster is defined as a serious disruption in the functioning of a community or society due to the interaction between a hazard event and the conditions of exposure, vulnerability, and capacity; resulting in human, material, economic, or environmental losses and impacts (Ge et al., 2020). According to (UNDRR, 2018a), three types of natural disasters – earthquakes, storms, and floods – have caused 1.12 million deaths, affected 2.85 billion people, and resulted in more than 2,647 billion US dollars in economic losses between 1998 and 2017. The threat of disasters is likely to increase as a result of global urbanization. Rapid and comprehensive damage assessment is crucial in the aftermath of a disaster to make informed and effective rescue decisions and minimize losses and impacts. Among these assessments, building damage assessment, aiming at providing information including the area and amount of damage, the rate of collapsed buildings, and the type of damage incurred by each building, is particularly critical in the early stages. This is because the distribution of damaged buildings is closely related to life-saving efforts in emergency response (Xie et al., 2016; Adriano et al., 2021). However, conducting field surveys after a disaster is difficult and dangerous, especially when transportation and communication systems are disrupted, impeding efficient on-site assessments. Earth observation (EO) technology, also known as remote sensing, provides a safe and efficient way to obtain building damage information over disaster areas due to its wide field-of-view, low cost, and contactless operation.
Two primary types of EO technology, optical and synthetic aperture radar (SAR), are commonly used for assessing building damage after disasters. Optical EO imagery has been the primary source and has been extensively studied for building damage assessment, because of its intuitive and easy-to-interpret nature. Initially, moderate-resolution optical EO data, exemplified by the Landsat series and Sentinel-2, was used for assessing building damage (Yusuf et al., 2001; Fan et al., 2019; Sandhini Putri et al., 2022). However, these were limited by spatial resolution and only provided broad approximations of affected areas, lacking accuracy for specific buildings, which is crucial for timely rescue. The new generation of very-high-resolution (VHR) optical sensors, such as IKONOS and WorldView, provide EO data with spatial resolutions at the meter and even sub-meter level, enabling finer assessments at the level of individual buildings (Freire et al., 2014). These data have been successfully used for building damage assessment after multiple natural disasters (Yamazaki and Matsuoka, 2007; Tong et al., 2012; Freire et al., 2014). While accurate building damage proxy maps can be obtained through visual interpretation by experts, this process is time-consuming, requires trained professionals, and is in practice too labor-intensive for rapid, large-scale assessments. Therefore, much of the recent literature has focused on developing automated methods for rapid building damage proxy mapping (Tong et al., 2012; Xie et al., 2016; Gupta et al., 2019; Zheng et al., 2021). Among these, artificial intelligence (AI), especially deep learning (DL) techniques, have proved particularly effective, significantly improving the efficiency and accuracy of building damage assessments. Early works were often specific to a single disaster event, with labels annotated for a particular disaster area to train the model, which was then used to generate building damage proxy maps for the same event (Xie et al., 2016; Xia et al., 2023). Since the associated training sets are limited to a narrow range of building types, damage patterns, and background land-cover distributions, the resulting models lack generalizability and struggle to produce accurate building damage proxy maps for new disaster events, which limits their practical use. The release of large-scale benchmark datasets including a variety of disaster scenarios and damage types, such as the xBD dataset (Gupta et al., 2019), opens up the possibility of using DL models to quickly and accurately map building damages after a newly occurred, previously unseen disaster. A number of state-of-the-art methods based on various DL models have been designed (Zheng et al., 2021; Shen et al., 2022; Guo et al., 2024; Wang et al., 2024; Chen et al., 2022; Kaur et al., 2023; Chen et al., 2024). For example, Zheng et al. (2021) trained DL models on the xBD dataset and applied them to mapping building damages on two unseen man-made disaster events. These studies showcase the power of employing DL models for building damage mapping. However, optical EO, being a passive sensing technique, needs suitable solar illumination and cloud-free weather conditions, which severely limits its application as an emergency tool for all-weather disaster response (Adriano et al., 2021). Actual disaster events, especially wildfires, floods, and storms, are often accompanied by less-than-ideal imaging conditions. For instance, Figure 1 shows EO data captured for a wildfire event that occurred in August 2023 in Hawaii, USA. The post-event optical image shown in Figure 1b does not provide clear surface information due to the effects of smoke from the wildfire. In contrast, SAR sensors employ active illumination with longer microwaves and can acquire images in adverse weather conditions, offering great potential for all-weather disaster response. As illustrated in Figure 1c, SAR imagery is unaffected by smoke and clearly shows buildings damaged by the wildfire.
@@
Due to the above advantages of SAR data and the development of associated processing technology, various SAR-based methods have been proposed for assessing building damage. These methods utilize intensity (Matsuoka and Yamazaki, 2005, 2010; Matsuoka et al., 2010), coherence (Yonezawa and Takeuchi, 2001; Arciniegas et al., 2007; Watanabe et al., 2016; Liu and Yamazaki, 2017), or polarization features (Yamaguchi, 2012; Chen and Sato, 2013; Watanabe et al., 2016; Karimzadeh and Mastuoka, 2017) for the building damage assessment at the block unit level, depending on the acquisition mode. Several studies have attempted to extend the block-level approach or have explored new approaches to move towards building instance level, using higher spatial resolution sensors like COSMO-SkyMed and TerraSAR-X (Liu et al., 2013; Brett and Guida, 2013; Chini et al., 2015; Ge et al., 2019). Some studies have also explored AI-based approaches for building damage assessment based on SAR data (Bai et al., 2018; Adriano et al., 2019; Bai et al., 2017; Li et al., 2023b). However, due to the lack of large-scale benchmark datasets (like xBD in the optical domain), these methods have focused on local regions and single disaster events, and their ability to generalize to other disaster events remains largely unknown. Challenges inherent to SAR data, such as oblique viewing angles, speckle noise, object occlusion, and geometric distortions, complicate the accurate mapping of buildings compared to optical imagery (Brunner et al., 2010; Adriano et al., 2021). Additionally, the limited availability of VHR SAR data also reduces its reliability as a source of pre-event data (Brunner et al., 2010). Considering these practical limitations, the most effective strategy for rapid, all-weather building damage assessment might arguably be to combine pre-event optical imagery, which provides accurate localization and detailed building information in the visible spectrum, with post-event SAR imagery, which captures structural information as a cue for building damages (Adriano et al., 2019). However, the image formation of SAR differs from that of an optical camera, both in terms of the physical properties of the imaged targets and in terms of the signal processing steps to create the image. Previous methods have attempted to align the two modalities with manually constructed statistical models (Stramondo et al., 2006; Chini et al., 2009; Brunner et al., 2010; Wang and Jin, 2012). However, these models are sensor-specific, requiring dedicated modeling for each sensor. DL offers a promising solution by automatically learning a high-dimensional feature space that aligns the two modalities. Still, a significant challenge remains: to train and evaluate a DL model, one must have access to a high-quality, large-scale dataset with comprehensive coverage of various disaster events and sufficient geographic diversity.
@@
To support AI research aimed at all-weather building damage mapping, we present BRIGHT, the first open, globally distributed, multimodal VHR dataset for building damage assessment. Advances in EO technology have enabled data providers like Capella and Umbra to offer VHR SAR imagery at resolutions higher than 1 meter per pixel. Benefiting from that progress, BRIGHT incorporates both pre-event optical imagery and post-event SAR imagery with spatial resolutions ranging from 0.3 meters to 1 meter per pixel. Such high resolution allows for detailed assessments at the individual building level, as required by emergency responders to guide targeted and effective rescue operations. Different event types are considered in BRIGHT: earthquakes, storms (hurricane, cyclone), wildfires, floods, and volcanic eruptions. These natural disasters account for 84 of the fatalities and 94 of the economic losses from 1998 to 2017 (UNDRR, 2018a). In addition to natural disasters, the BRIGHT dataset further considers disasters caused by human activity, such as accidental explosions or armed conflicts, which also pose significant threats to human life and infrastructure, can occur unexpectedly and require a rapid response (UNDRR, 2018b). The inclusion of these events ensures the dataset is comprehensive and can be used to develop robust AI models that are effective across a range of disaster scenarios. The events in BRIGHT cover 12 different regions distributed around the globe, with a focus on developing countries where external assistance is most urgently needed after a disaster. BRIGHT’s labels are multi-level annotations that distinguish between damaged and entirely destroyed buildings.
@@
(Nguyen et al., 2017) | Yes | Images on social media | N/A | 1 | 3 natural disasters | N/A | Image-level |
@@
(Sun et al., 2024) | Yes | Optical and SAR EO | <1 | 1 | Earthquake | 4,029 | Pixel-level |
@@
BDD (Adriano et al., 2021) | No | optical and SAR EO | 1*-10 | 9 | 3 natural disasters | 123,453 | Pixel-level |
BRIGHT | Yes | optical and SAR EO | 0.3-1 | 12 + N† | 5 natural disasters human-made disasters | >350,000 | Pixel-level |
@@ 0.1 Comparison with existing datasets
The comparison between BRIGHT and existing datasets for building damage assessment is summarized in Table 1. Most current building damage assessment datasets are limited in scale and scope due to the scarcity of disaster events, the limited availability of corresponding open-source EO data and the annotation effort (Rahnemoonfar et al., 2021; Gupta and Shah, 2021; Kaur et al., 2023). Due to the high cost and time required for pixel-level labeling, some datasets, such as those provided by (Fujita et al., 2017; Nguyen et al., 2017; Cheng et al., 2021; Xue et al., 2024), offer image-level labeling, indicating only whether an image contains damaged buildings. Although useful, this information lacks the spatial precision needed to guide specific rescue operations. The xBD dataset (Gupta et al., 2019) is currently the largest open data collection, covering six natural disasters in 15 regions with more than 700,000 building instances. However, xBD includes only optical EO data and therefore does not support all-weather disaster response. Sun et al. (2024) have created a multimodal dataset, but it is limited to a single disaster event and contains only about 4,000 building instances. The small size makes it challenging to train DL models, and significantly limits the transferability of the trained models. The dataset most similar to our work is BDD, proposed by Adriano et al. (2021). The main differences between BDD and BRIGHT are: First, BRIGHT covers more disaster events and building instances, including both natural and human-made disasters. Second, BRIGHT has higher spatial resolution; while the highest resolution of SAR images in BDD is 1 meter, BRIGHT provides finer detail with spatial resolution from 0.3 meter to 1 meter, allowing for the detection of subtle structural damage in individual buildings. Third, and perhaps most important, the re-distribution of BDD is restricted, whereas BRIGHT is an open-source dataset publicly available to the global community. Beyond the ones listed in Table 1, there are further datasets targeted at monitoring hazardous events related to disasters, but not concerned with building damage assessment, including landslides (Ghorbanzadeh et al., 2022; Meena et al., 2023), floods (Bonafilia et al., 2020; Zhang et al., 2023) and wildfires (Artés et al., 2019; Huot et al., 2022; He et al., 2024).
@@ 0.2 Main contribution
The contributions of this paper are twofold: 1) We present BRIGHT, the first multimodal building damage dataset with a spatial resolution greater than 1 meter, and make it publicly available to the community. BRIGHT employs a combination of pre-event optical imagery and post-event SAR imagery with abundant disaster events and rich geographic diversity, thereby supporting the study of AI-based multimodal building damage mapping, especially in developing countries; 2) We evaluate a series of contemporary DL baselines on BRIGHT. The results of these experiments, along with the source code and model weights, are also made publicly available in support of future research and a benchmark for the current potential of EO-based disaster response. The trained models may serve as a basis for methodological advances or to build damage assessment pipelines for real-world applications.

### v2

@@ Abstract
Disaster events occur around the world and cause significant damage to human life and property. Earth observation (EO) data enables rapid and comprehensive building damage assessment, an essential capability in the aftermath of a disaster to reduce human casualties and inform disaster relief efforts. Recent research focuses on the development of artificial intelligence (AI) models to achieve accurate mapping of unseen disaster events, mostly using optical EO data. However, solutions based on optical data are limited to clear skies and daylight hours, preventing a prompt response to disasters. Integrating multimodal EO data, particularly the combination of optical and synthetic aperture radar (SAR) imagery, makes it possible to provide all-weather, day-and-night disaster responses. Despite this potential, the development of robust multimodal AI models has been constrained by the lack of suitable benchmark datasets. In this paper, we present a Building damage assessment dataset using veRy-hIGH-resoluTion optical and SAR imagery (BRIGHT) to support AI-based all-weather disaster response. To the best of our knowledge, BRIGHT is the first open-access, globally distributed, event-diverse multimodal dataset specifically curated to support AI-based disaster response. It covers five types of natural disasters and two types of man-made disasters across 14 regions worldwide, with a particular focus on developing countries where external assistance is most needed. The optical and SAR imagery in BRIGHT, with a spatial resolution between 0.3 and 1 meters, provides detailed representations of individual buildings, making it ideal for precise damage assessment. In the experiments, we have tested seven advanced AI models trained with BRIGHT to validate the transferability and robustness. Beyond that, BRIGHT also serves as a challenging benchmark for a variety of tasks under disaster scenarios, including unsupervised domain adaptation (UDA), semi-supervised learning (SSL), unsupervised multimodal change detection (UMCD), and unsupervised multimodal image matching (UMIM). The experimental results serve as baselines to inspire future research and model development. The dataset, along with code and trained models, is available at https://github.com/ChenHongruixuan/BRIGHT and will be updated as new disaster data becomes available. BRIGHT also serves as the official dataset for the 2025 IEEE GRSS Data Fusion Contest Track II. We hope that this effort will promote the development of AI-driven EO methods in support of people in disaster-affected areas.
@@
A disaster is defined as a severe disruption in the functioning of a community or society due to the interaction between a hazard event and the conditions of exposure, vulnerability and capacity resulting in human, material, economic or environmental losses and impacts (Ge et al., 2020). According to the United Nations Office for Disaster Risk Reduction (UNDRR), between 1998 and 2017, natural disasters such as earthquakes, storms, and floods have affected 2.85 billion people and caused approximately 1.12 million deaths. These disasters have also resulted in economic losses of US$2,647 billion in disaster-affected countries (UNDRR, 2018a). The threat of disasters is likely to increase due to global urbanization (Kreibich et al., 2022; Bastos Moroz and Thieken, 2024). Rapid and comprehensive damage assessment is crucial in the aftermath of a disaster to make informed and effective rescue decisions to minimize losses and impacts. Building damage assessment aims to provide information, including the area and amount of damage, the rate of collapsed buildings, and the type of damage each building has incurred. This information is critical in the early stages of a disaster, as the distribution of damaged buildings is closely related to life-saving efforts in an emergency response (Xie et al., 2016; Adriano et al., 2021). Conducting field surveys after a disaster is difficult and dangerous, especially when transportation and communication systems are disrupted, preventing efficient on-site assessments. Earth observation (EO), also known as remote sensing, provides a safe and efficient way to obtain information on building damage in disaster areas because of its wide field of view and contactless operation.
The EO technologies commonly used for assessing building damage after disasters are optical and synthetic aperture radar (SAR). Optical imagery is a primary source for building damage assessment because of its intuitive and easy-to-interpret nature. For example, moderate-resolution optical EO data from the Landsat series and Sentinel-2 have been used to assess building damage (Yusuf et al., 2001; Fan et al., 2019; Sandhini Putri et al., 2022). Landsat and Sentinel-2 data are limited in spatial resolution and only provide broad approximations of affected areas, which lack accuracy for specific buildings crucial for timely rescue. The new generation of very-high-resolution (VHR) optical sensors, such as IKONOS and WorldView, provide EO data with spatial resolutions at a meter and a sub-meter level, enabling finer assessments at the level of individual buildings (Freire et al., 2014). These data have been used successfully for the assessment of building damage (Yamazaki and Matsuoka, 2007; Tong et al., 2012; Freire et al., 2014).
While accurate building damage maps can be obtained by human-expert visual interpretation of optical images, this process is time-consuming and labor-intensive for rapid large-scale assessments. In addition, it requires trained professionals. Therefore, recent literature has focused on developing automated methods for rapid building damage mapping (Tong et al., 2012; Xie et al., 2016; Gupta et al., 2019; Zheng et al., 2021). Among these, deep learning (DL) techniques have significantly improved the efficiency and accuracy of building damage assessment. Earlier work was toward a single disaster event with labels annotated for a particular disaster area to train a model, which is then used to generate building damage maps for the same event (Xie et al., 2016; Xia et al., 2023). However, since training data were limited to a few building types, damage patterns, and background land cover distributions, the resulting models mostly lack generalizability and struggle to produce accurate building damage maps for new disaster events, which limits their practical use. Recent large-scale benchmark datasets, for example, the xBD dataset (Gupta et al., 2019) of different types of disaster scenarios and damages, have made it possible to adopt DL models to quickly and accurately map building damages after a newly occurred, previously unseen disaster (Zheng et al., 2021; Shen et al., 2022; Guo et al., 2024; Wang et al., 2024; Chen et al., 2022a; Kaur et al., 2023; Chen et al., 2024). For example, Zheng et al. (2021) trained DL models on the xBD dataset and applied them to map the damages of buildings in two unseen human-made disaster events. These studies have demonstrated the effectiveness of DL models for building damage mapping.
@@
The optical EO technology uses a passive sensing technique, which requires solar illumination and cloud-free weather conditions. This severely limits the application of optical images in an emergency tool for all-weather disaster response (Adriano et al., 2021). In contrast, SAR sensors employ active illumination with longer microwaves and can acquire images in adverse weather conditions, offering great potential for all-weather disaster response. Most disaster events, especially wildfires, floods, and storms, are often accompanied by less-than-ideal imaging conditions. For example, Figure 1 shows EO imagery captured for a wildfire event that occurred in August 2023 in Hawaii, USA. The post-event optical image shown in Figure 1b does not provide clear surface information due to the effects of the wildfire smoke. However, the SAR image illustrated in Figure 1c is not affected by smoke and clearly shows the buildings damaged by the wildfire. Due to the advantages of SAR imagery, various SAR-based methods have been proposed for assessing building damage. These methods utilize intensity (Matsuoka and Yamazaki, 2005, 2010; Matsuoka et al., 2010), coherence (Yonezawa and Takeuchi, 2001; Arciniegas et al., 2007; Watanabe et al., 2016; Liu and Yamazaki, 2017), and polarization features (Yamaguchi, 2012; Chen and Sato, 2013; Watanabe et al., 2016; Karimzadeh and Mastuoka, 2017) to assess building damage at a block unit level, depending on the acquisition mode. Several studies have attempted to extend the block-level approach and have explored new approaches at the building instance level using higher spatial resolution sensors such as COSMO-SkyMed and TerraSAR-X (Liu et al., 2013; Brett and Guida, 2013; Chini et al., 2015; Ge et al., 2019). DL-based methods have also been explored with SAR data to assess building damage (Bai et al., 2018; Adriano et al., 2019; Bai et al., 2017; Li et al., 2023b). However, because of the lack of large-scale benchmark datasets, such as xBD in the optical domain, these methods have focused on local regions and single disaster events, and their ability to generalize to other disaster events remains largely unknown. Inherent challenges of SAR data, such as oblique viewing angles, speckle noise, object occlusion, and geometric distortions, complicate accurate building damage mapping compared to optical imagery (Brunner et al., 2010; Adriano et al., 2021). Furthermore, the limited availability of the VHR SAR data reduces its reliability as a source of pre-event data (Brunner et al., 2010). Considering these practical limitations, the most effective strategy for rapid assessment of building damage in all weather could arguably be to combine pre-event optical images, which provide accurate localization and detailed building information in the visible spectrum, with post-event SAR images, which capture structural information as a cue for building damage (Adriano et al., 2019). Previous methods have attempted to align the two modalities with traditional statistical models (Stramondo et al., 2006; Chini et al., 2009; Brunner et al., 2010; Wang and Jin, 2012). These statistical models are sensor-specific and require dedicated modeling for each sensor. DL methods offer a promising solution by automatically learning a high-dimensional feature space that aligns the two modalities. However, to train a DL model, one must have access to a high-quality large-scale dataset with comprehensive coverage of various disaster events and sufficient geographic diversity. This is a significant challenge that remains and needs to be addressed.
@@
To support DL research aimed at all-weather building damage mapping, we present BRIGHT, the first open and globally distributed multimodal VHR dataset for building damage assessment. Advances in EO technology have enabled data providers like Capella111https://www.capellaspace.com/ and Umbra222https://umbra.space/ to offer VHR SAR imagery at a sub-meter level resolution per pixel. This allows for detailed building damage assessments at the individual building level to guide targeted and effective rescue operations as required by emergency responders. Benefiting from the progress that has been made in EO, BRIGHT incorporates both pre-event optical imagery and post-event SAR imagery with spatial resolutions ranging from 0.3 meters to 1 meter per pixel. The types of disaster events considered in BRIGHT are earthquakes, storms (e.g., hurricane, cyclone), wildfires, floods, and volcanic eruptions. These natural disasters accounted for 84 of the fatalities and 94 of the economic losses between 1998 and 2017 (UNDRR, 2018a). In addition to natural disasters, the BRIGHT dataset further considers disasters caused by human activity, such as accidental explosions and armed conflicts, which also pose significant threats to human life and infrastructure and can occur unexpectedly, requiring a rapid response (UNDRR, 2018b). The events cover 14 different regions distributed around the globe with a focus on developing countries where external assistance is most urgently needed after a disaster. The labels are manually annotated with multi-level annotations that distinguish between damaged buildings and completely destroyed buildings.
@@
| (Nguyen et al., 2017) | Yes | Images on social media | N/A | 4 | 3 natural disasters | N/A | Image-level |
@@
| QQB (Sun et al., 2024b) | Yes | Optical and SAR EO | <1 | 1 | Earthquake | 4,029 | Pixel-level |
@@
| BDD (Adriano et al., 2021) | No | Optical and SAR EO | 1.2-3.3 | 9 | 3 natural disasters | 123,453 | Pixel-level |
| BRIGHT | Yes | Optical and SAR EO | 0.3-1 | 14 | 5 natural disasters 2 human-made disasters | 384,596 | Pixel-level |
@@ 0.1 Comparison with existing datasets
The comparison between BRIGHT and existing datasets for building damage assessment is summarized in Table 1. Most current building damage assessment datasets are limited in scale and scope due to limited available disaster events with corresponding open-source EO data and annotation effort (Rahnemoonfar et al., 2021; Gupta and Shah, 2021; Kaur et al., 2023). Because of the high cost and time required for pixel-level labeling, some of the existing datasets provide image-level labeling, indicating only whether an image contains damaged buildings (Fujita et al., 2017; Nguyen et al., 2017; Cheng et al., 2021; Xue et al., 2024). Although these image-level labeling datasets have served the community well, they lack the spatial precision needed to guide specific rescue operations. The xBD dataset (Gupta et al., 2019) is currently the largest open data collection, covering six natural disasters in 15 regions with more than 700,000 building instances. The xBD includes only optical EO data. Therefore, it does not support all-weather disaster response. Sun et al. (2024b) introduced a multimodal dataset, but it is limited to a single disaster event and contains only about 4,000 building instances. The small size makes it challenging to train DL models and limits the transferability of the trained models. The dataset most similar to BRIGHT is the BDD proposed by Adriano et al. (2021). The main differences between BDD and BRIGHT datasets are: (1) BRIGHT covers more disaster events and building instances, including both natural and human-made disasters. (2) BRIGHT has higher spatial resolution; while the highest resolution of SAR images in BDD is 1.2 meters, BRIGHT provides finer detail with spatial resolution from 0.3 meters to 1 meter to enable the detection of subtle structural damage in individual buildings. (3) Perhaps the most important difference is that whereas the re-distribution of BDD is restricted, BRIGHT is an open-source dataset publicly available to the global community. Apart from the datasets listed in Table 1, there are other datasets targeted at monitoring hazardous events related to disasters, including landslides (Ghorbanzadeh et al., 2022; Meena et al., 2023), floods (Bonafilia et al., 2020; Zhang et al., 2023) and wildfires (Artés et al., 2019; Huot et al., 2022; He et al., 2024), but are not concerned with building damage assessment.
@@ 0.2 Main contribution
The contributions of this paper are twofold:
1.
We present BRIGHT, the first multimodal building damage dataset with a spatial resolution greater than 1 meter, which is publicly available to the community. BRIGHT employs a combination of pre-event optical imagery and post-event SAR imagery with various disaster events and rich geographic diversity to support the study of AI-based multimodal building damage mapping, especially in developing countries.
2.
We evaluate a suite of contemporary models on BRIGHT to establish robust baselines. Beyond supervised deep learning, BRIGHT can support a wide range of AI-based methods. It enables research in UDA, SSL, UMCD, and UMIM, among others. To demonstrate its utility, we benchmark a suite of representative models across several of these tasks. All experimental results, along with source code and pretrained model weights, are publicly released to provide strong baselines and accelerate future developments in EO-based disaster response.